how to solve this (1/3)3 divided by 1/2 (-9/4)?
Answer by user_dude2008(1862)

(1/3)3 = 3/3 = 1

(1/2)(-9/4) = -9/8



(1/3)3 divided by 1/2 (-9/4) = 1/(-9/8) = -8/9

Answer: -8/9